Solve by using the perfect squares method.
x2 + 4x + 4 = 0


I am not quite sure how to do this problem.

Found 2 solutions by checkley77, solver91311:
Answer by checkley77(12569)

x^2+4x+4=0
x^2+4x+4=4-4
x^2+4x+4=0
(x+2)^2=0

Answer by solver91311(16877)





Since 2 + 2 = 4 and 2 X 2 = 4, the left had expression is a perfect square trinomial:



Using the Zero Product Rule



or



John